A satellite orbits the earth at a speed of 8*10∆3m/s.calculate the period of orbit of the satellite if the distance between the centre of the earth and satellite is 6400km

1
Expert's answer
2022-03-20T18:48:18-0400

The orbital period of rotation of the satellite

"T=\frac{2\pi R}{v}\\
=\rm\frac{2\times 3.14\times 6400\times 10^3\: m}{8\times 10^3\: m/s}=5024\: s=83.7\: min"